# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 6 In State Forward:
66       Initialize $W_{time}=0$ and change to state Gather2.
@@ 79             Move in a clockwise direction with all the pebbles at the $Current-Node$
until it satisfies the condition, where it finds a node on which a total of 3 agents exists. After which they change their state to Coloc.
@@
5.1.1 Correctness and Complexity
In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Scat-Pbl.
@@ 5.1.1 Correctness and Complexity
PerpExplore-Scat-Pbl guarantees to solve PerpExploration-BBh problem if no agent is consumed by the byzantine black hole.
Proof 5.2.
Let us define a segment for an agent $a_{1}$ be $Seg(a_{1})$, where $Seg(a_{1})$ is defined to be the collection of consecutive nodes in the clockwise direction, starting from its own $home$ and its clockwise nearest $home$ (which is the $home$ of another agent say, $a_{2}$). Note that, since no agent is consumed by the black hole, any agent, say $a_{1}$, while executing the Algorithm 3, starts moving in clockwise direction in state Forward, reaches the nearest clockwise $home$, i.e., the last node of this segment $Seg(a_{1})$, then changes its state to Wait1, and after waiting for $n-size$ rounds, again changes its state to Fetch. In state Fetch, the agent basically moves counter-clockwise, until and unless it reaches its own $home$, which is nothing but the first node of this segment $Seg(a_{1})$. Also, note that, each agent, changes their state to Forward and Fetch together at the same round, while exploring their respective segments. Now, note that, the union of all these segments is basically the whole ring $R$ (i.e., more specifically $\cup^{4}_{i=1}Seg(a_{i})=R$) . Hence, this guarantees that the ring is perpetually explored if no agent is consumed by the black hole.
@@
If an agent is consumed by the byzantine black hole, then within an additional finite rounds, there exists exactly one agent, which changes its state to Gather1.
Proof 5.4.
Let $a_{1}$ be an agent exploring $Seg(a_{1})$. Without loss of generality, let the black hole be in $Seg(a_{1})$. Also let $a_{2}$ be another agent which is exploring the segment $Seg(a_{2})$, where $home$ of $a_{2}$ is the common node between $Seg(a_{1})$ and $Seg(a_{2})$. Also, let $a_{3}$ be an agent which is exploring the segment $Seg(a_{3})$, where the $home$ of $a_{3}$ is the common node between $Seg(a_{2})$ and $Seg(a_{3})$. Note that, $a_{1}$ can be consumed by the black hole, either in state Forward or in state Fetch.
  • If $a_{1}$ is consumed while in state Forward, then it fails to bring back the pebble from the $home$ of $a_{2}$, while it is scheduled to change its state to Fetch, after ending the states Forward and Wait1. Note that, since we have assumed that the black hole is in $Seg(a_{1})$, and $Seg(a_{1})\cap Seg(a_{2})=\{home\text{ of }a_{2}$}. Thus, $a_{2}$ must bring a pebble from the $home$ of $a_{3}$ to it’s own $home$, after it changes its state to Fetch. So, now as $a_{2}$ reaches its own $home$ along with the pebble it carried from the $home$ of $a_{3}$, it finds that there already exists another pebble, which the agent $a_{1}$ is unable to carry back to its own $home$, as it is already consumed by the black hole. This causes an anomaly for $a_{2}$ at its own $home$ (i.e., encountering more than one pebble). After detecting the anomaly $a_{2}$ immediately changes its state to Gather1. Note that, after $a_{1}$ is consumed by the black hole, it has taken an additional at most $size-1+(n-size)+1+size+1=n+size+1$ 333calculates the number of rounds it took $a_{2}$ to change to state Gather1, where at most $size-1$ rounds required in state Forward after $a_{1}$ gets consumed by the black hole, $(n-size)$ rounds for changing to state Wait1 and waiting, 1 round for changing from Wait1 to Fetch, then $size$ rounds for moving in state Fetch and returning back to its $home$, and lastly 1 round for changing to state Gather1 from Fetch. rounds to change the state of $a_{2}$ to Gather1. Further in this case we show that no other agents except $a_{2}$ changes to state Gather1. Let $t^{\prime}$ be the round when $a_{2}$ changes its state to Gather 1. And $t<t^{\prime}$ be the last round before $t^{\prime}$ when all the agents moved into state Forward. Then $t^{\prime}=t+n+1+size+1$. Next $a_{2}$ waits for $n-size-1$ rounds that is until round $t+2n+1$ and starts moving counter clockwise being in state Gather1. Note that by observation 1, the other agents stay at their home until round $t+3n+1$. So from $t+2n+2$ until $t+3n+1$, these $n-1$ rounds are enough for $a_{2}$ to meet the other two while they are still being in the state Wait2. So they change their state to Gather2 and eventually to Coloc. As an agent can not change state to Gather1 from state Coloc or Gather2, no other agent changes to state Gather1.
  • If $a_{1}$ is consumed while in state Fetch. In this case, $a_{2}$ after returning back to its $home$ carrying the pebble from the nearest clockwise $home$, i.e., the $home$ of $a_{3}$, finds no anomaly. It is because, $a_{1}$ has been consumed by the black hole while in state Fetch, so that means, the pebble at $home$ of $a_{2}$ has been carried by $a_{1}$. Since, $a_{2}$ does not encounter any anomaly, hence it drops the pebble at its $home$, changes to state Wait2, waits at $home$ for $12n-size-$ 1 rounds and then changes to state Forward again. So, while executing the same execution, after reaching the clockwise nearest $home$, i.e., $home$ of $a_{3}$, it again changes its state to Wait1, waits for $n-size-1$ rounds, and further changes its state to Fetch. Note that, here at the $home$ of $a_{3}$, there exists a pebble which was earlier dropped by $a_{3}$, while it changed its state from Fetch to Wait2. So, $a_{2}$ carries this pebble and reaches its own $home$, where it finds the anomaly. It is because, as $a_{1}$ has been consumed by the black hole, it is unable to carry the pebble dropped by $a_{2}$ during its last state change from Fetch to Wait2. This phenomenon triggers $a_{2}$ to change its state to Gather1. Note that, since $a_{1}$ has been consumed by the black hole, it takes an additional $(size-1)+1+(2n-size)+size+1+(n-size)+size+1=3n+size+2$ 222calculates the number of rounds it took $a_{2}$ to change to state Gather1, where $size-1$ rounds required in state Fetch after $a_{1}$ gets consumed by the black hole, 1 round for changing its state from Fetch to Wait2, after this the agent waits for $12n-size-$ 1 rounds and in one additional round changes to state Forward, in this state it takes $size$ rounds, 1 round required to change to state Wait1, then $n-size-1$ rounds the agent waits, whereas 1 more round required to change to state Fetch, finally $size$ rounds in state Fetch, and lastly 1 round for changing its state to Gather1rounds to change $a_{2}$’s state to Gather1. And as shown for the first case it can be similarly showed that except $a_{2}$ no other agent changes state to Gather1
Lemma 5.5.
After the first agent changes its state to Gather1, it takes an additional at most $22n-$ 2 rounds, for the remaining agents to gather at a single node (i.e., the nearest counter-clockwise $home$ of the black hole node) along with their pebbles.
Proof 5.6.
Let $a_{2}$ be the agent which changes to state Gather1 at $t$-th round (where $t>0$), due to the fact that it encounters an anomaly at its $home$ because $a_{1}$ (w.l.o.g) has been consumed by the black hole, and the length of the segment it explores be denoted by $size$ (i.e., length of $Seg(a_{2})=size$). Also, let $t_{0}$ ($0<t_{0}<t$) be the last round when each agent changed their states to Forward. Note that $t$ should be exactly equal to $t_{0}+size+1+(n-size-1)+1+size+1=t_{0}+n+size+2$ 333movement in state Forward for $size$ rounds, then one round for change to state Wait1, then stay in state Wait1 for $n-size-1$ rounds, after which change to state Fetch for one round, and move in state Fetch for $size$ rounds, and finally change to state Gather1 for one round. Note that $a_{2}$ after changing its state to Gather1 (current round is $t_{0}+n+size+2$, which is also the $t-$th round according to our definition of $t$), further waits for $n-size-1(\leq n-2$, as $size\geq 1$ ) rounds (current round $t_{0}+2n+1$), and starts moving clockwise from $(t_{0}+2n+2)-$th round. On the other hand, an alive agent which was in state Fetch at $t-$th round, must have changed its own state to Fetch at $(t_{0}+n+1)-$th round (refer Observation 5.1). From Fetch, it takes further at most $n$ (since length of a segment can be at most $n-1$) rounds to change to state Wait2. Note that any other alive agent except $a_{2}$ must change to state Wait2 after completion of state Fetch as by Lemma 5.3, exactly one agent (here $a_{2}$) already changed its state to Gather1 from Fetch at round $t=t_{0}+n+size+2$). So, any agent except $a_{2}$ at the round $(t_{0}+2n+1)-$th round must be in state Wait2, and should remain in this state until $(t_{0}+3n+2)-$th round, if no anomaly is detected (refer Observation 1) or meets with other agents. So, any alive agent except $a_{2}$ remains at their $home$ within $t_{0}+2n+1$ to $t_{0}+3n+2$ rounds. This is by Lemma 5.3, only $a_{2}$ being the first agent to detect anomaly can change its state to Gather1, whereas the remaining agent can only change its state to Gather2 or Coloc, when they meet with $a_{2}$, while in state Wait2. Further, note that an alive agent except $a_{2}$, waits at their respective $home$ for $n+1$ rounds since, $a_{2}$ starts moving clockwise being in state Gather1 (after $(t_{0}+2n+1)-$th round). Now, when $a_{2}$ reaches its clockwise neighbor (say) $a_{3}$, at its $home$. $a_{3}$ must be at state Wait2, in this case, $a_{3}$ changes its state to Gather2, and starts moving clockwise along with $a_{2}$ carrying their respective pebbles, until they meet another agent, (say) $a_{4}$, at its $home$, while in state Wait2. After which they all change to state Coloc, which also guarantees that they have at least 3 pebbles at the current node (since $a_{2}$ carried at least 2 pebbles from its own $home$ where it detects the anomaly, $a_{3}$ also carried its one pebble present at its own $home$). Now since no agent can change to state Gather1 from Gather2 or Coloc, $a_{2}$ is the only agent that changes to state Gather1. Note that, before gathering $a_{2}$, $a_{3}$ and $a_{4}$ can never be destroyed, as they traverse only the nodes in $\cup^{4}_{i=2}Seg(a_{i})$, whereas the black hole is in $Seg(a_{1})\backslash\{home$ of $a_{1}$, $home$ of $a_{2}\}$.
@@
In the discussion of our Algorithm 3, we have considered that 4 agents are scattered along 4 distinct nodes (i.e., each node with multiplicity 1). Here we describe how our algorithm (Algorithm 3) works for the remaining cases, (i.e., when 4 agents are scattered among 3 or 2 nodes initially) by slight modification. Observe that, in these remaining cases there exists at least one node, where multiplicity is greater than 1 initially. Let there exists a node with multiplicity $3$, in this case, these agents directly start executing PerpExplore-Coloc-Pbl, while execution if they encounter the fourth agent somewhere along $R$, they just ignore this agent (note that the IDs, of all the 3 initially co-located agents are already collected by each other) while executing their current algorithm. On the other hand, if initially there exists a node with multiplicity greater than 1 and less than 3, then in that case, only the lowest ID agent (say $a_{1}$) at the current node, changes its state Forward, whereas the other agent at the current node (say $a_{2}$) changes its state Backup-Wait at some round say $t_{0}$ $(>0)$. The agent in state Forward, continue executing the algorithm PerpExplore-Scat-Pbl, whereas the agent in state Backup-Wait, waits at the current node for $12n+$ 1 rounds (i.e., until $t_{0}+2n+1$-th round) and then checks for anomaly in the next round (i.e., if the current node has more pebble than current number of agents). If such anomaly exists after $t_{0}+2n+1$ round, then that implies that $a_{1}$ is already in state Gather1. It is because, $a_{1}$ while it returned back to its $home$ carrying a pebble in state Fetch at round $t_{0}+n+size+2$ round (where $size$ is the length of $Seg(a_{1})$), encounters the first anomaly, i.e., the number of pebble is more than number agents. In that case, it directly changes its state to Gather1, and further waits at $home$ for $n-size-1$ rounds, i.e., till $t_{0}+2n+1$ rounds. After which it starts to move clockwise, and on the other hand $a_{2}$ also starts moving clockwise, while changing its state to Gather2 from Backup-Wait. This guarantees that if the anomaly is detected at a multiplicity then both $a_{1}$ and $a_{2}$ moves together to gather, along with the third agent which is in state Wait2 (as described in Lemma 5.5) and changes its state to Gather2 whenever it finds #$\#Agent$ at its $home$ is more than one. Now if the anomaly is not detected at the multiplicity, that is $a_{2}$ at round $t_{0}+2n+1$ finds no anomaly then it moves into state Backup whereas $a_{1}$ is in state Wait2 at the same node. Both of them waits until round $13n+$ 1 and at round $23n+$ 2, $a_{1}$ changes its state to Forward again (refer to Observation 1) and $a_{2}$ changes its state to Backup-Wait again. If some other agent detects anomaly it must meet $a_{1}$ and $a_{2}$ at their $home$ at some round $t$ where $t_{0}+2n+2\leq t\leq t_{0}+3n+1$. In that case they find that there are 3 agents at their $home$ and all of them change to state Coloc and executes algorithm PerpExplore-Coloc-Pbl.
Next we can have the following theorem
Theorem 5.8.
A set of 4 synchronous scattered agents with one pebble each is necessary and sufficient to solve the PerpExploration-BBH problem on a ring $R$, when the initial starting locations of the agents are different along $R$.
5.2 Communication: Whiteboard

### v2

@@ 6 In State Forward:
66       Initialize $W_{time}=0$ and change to state Gather2 and moves clockwise.
@@ 79             Move in a clockwise direction with all the pebbles at the $Current-Node$
An agent in state Gather1, is the first agent to detect the anomaly, hence it is first instructed to wait at the current node, i.e., its $home$ for the first $n-size-1$ rounds, and then start moving in a clockwise direction, while accompanying all the pebbles at the current node  [^2] Note that we have considered that in this case an agent, while in state Gather1 or Gather2 can carry more than one pebble, which can be easily restricted to one agent can carry only one pebble, but in that case the agent needs to return and carry one pebble each time and reach back to its current node.
This movement in the clockwise direction continues until the number of agent at the current node is exactly 3, after which it changes to state Coloc.
An agent in state Gather2 (i.e., the alive agents which are not the first to detect the anomaly), first checks whether the current node has exactly 3 agents on it, if not then it moves in a counter-clockwise direction, while accompanying the pebbles present at its $home$, until it satisfies the condition, where it finds a node on which a total of 3 agents exists. After which they change their state to Coloc.
@@
10.2 Correctness and Complexity
In this section, we discuss the correctness and complexity of our algorithm PerpExplore-Scat-Pbl. Before that, let us define an iteration of the algorithm PerpExplore-Scat-Pbl. An iteration is a set of $13n+$ 1 consecutive rounds starting from the round in which all alive agents change to state Forward together. Next we prove the following lemma.
@@ 10.2 Correctness and Complexity
If no agents are destroyed by the Byzantine black hole until the completion of $p$-th iteration, then for each $j-$th iteration, every node of the ring $R$ is visited by at least one agent, where $j\in\mathbb{N}^{+},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ j\leq p$.
Proof 10.3.
Let $a_{0},a_{1},a_{2}$ and $a_{3}$ be 4 agents initially located at $h_{0},h_{1},h_{2}$ and $h_{3}$ respectively where, $h_{0},h_{1},h_{2},h_{3}$ are in clockwise order starting from $h_{0}$. By $Seg(a_{i})$ we denote the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$. Note that in any of the $j-$th iteration ($j\leq p$), an agent $a_{i}$ (for all $i\in\{0,1,2,3\}$) moves clockwise along $Seg(a_{i})$ starting from $h_{i}$ while in state Forward and reaches $h_{(i+1)\pmod{4}}$. Then $a_{i}$ in the same iteration moves back to $h_{i}$ again while being in state Fetch. Thus each node of $Seg(a_{i})$ is explored twice in $j$-th iteration by $a_{i}$, where $j\leq p$. Since for any $v\in R$, $v\in Seg(a_{j}),$ for some $j\in\{0,1,2,3\}$, and in each iteration $a_{j}$ explores each node of $Seg(a_{j})$ until $p-$th iteration, the result follows.
Corollary 10.4.
If no agents are destroyed, the agents can perpetually explore a ring execuing the algorithm PerpExplore-Scat-Pbl.
Next we prove, after the consumption of one agent by the Byzantine black hole, during a subsequent iteration or at the current iteration, at most one agent can change its state to Gather1 and gathers with two more agents before that iteration ends.
@@
In an iteration, at most one agent can change its state to Gather1 executing PerpExplore-Scat-Pbl. Furthermore, before the end of that iteration the agent gathers with two more agents at a safe node with at least 3 pebbles.
Proof 10.6.
Let $a_{0},a_{1},a_{2}$ and $a_{3}$ be 4 agents initially located at $h_{0},h_{1},h_{2}$ and $h_{3}$ respectively where, $h_{0},h_{1},h_{2},h_{3}$ are in clockwise order starting from $h_{0}$. By $Seg(a_{i})$ we denote the clockwise arc starting from $h_{i}$ and ending at $h_{(i+1)\pmod{4}}$. Without loss of generality, let the first agent that is destroyed by the Byzantine black hole be $a_{0}$ during the $p-$th iteration. So, the Byzantine black hole $v_{b}$, must be in $Seg(a_{0})$. Since, no agents changes to state Gather1, before an agent is destroyed. Hence, until $(p-1)-$th iteration, there does not exist any agent, which changes its state to Gather1. Now, we have following two cases, which are based on the state $a_{0}$ was in, before it gets destroyed by the Byzantine black hole.
Case-I: $a_{0}$ was in state Forward when it was destroyed. So, $a_{0}$ fails to reach $h_{1}$ to bring the pebble, left by $a_{1},$ back to $h_{0}$ during $p-$th iteration. So, when $a_{1}$ reaches $h_{1}$ at the $p-$th iteration, along with the pebble it collected from $h_{2}$, it finds two pebble at the node $h_{1}$. This happens at $(n+2+|Seg(a_{1})|)-$th round of $p-$th iteration. In the same round $a_{1}$ changes to state Gather1. Next it waits an additional $n-|Seg(a_{1})|-1$ rounds and then in the $(2n+2)-$ th round of the $p-$ th iteration, it starts moving clockwise along with both the pebbles. Note that, all other alive agents (i.e., $a_{2}$ and $a_{3}$) reaches their home by the $(2n+2)-$th round of the $p-$th iteration. Also upon reaching their respective $home$, $a_{2}$ and $a_{3}$ both sees exactly one pebble (i.e., the pebble they carried back to their $home$) so they change their state to Wait2. By Observation 1, from $(2n+2)$-th round of the iteration upto the end of the iteration (i.e., $(3n+1)-$th round of the iteration), both $a_{2}$ and $a_{3}$ stays at $h_{2}$ and $h_{3}$, respectively in state Wait2. So, after $a_{1}$ starts moving clockwise in the $(2n+2)-$th round, it first meets $a_{2}$ in round $2n+1+|Seg(a_{1})|$ (which is $\leq 3n-1$ as $|Seg(a_{1})|\leq n-2$) of the $p-$th iteration. And then meets $a_{3}$ at round $2n+1+|Seg(a_{1})|+|Seg(a_{2})|\leq 3n$ (as $|Seg(a_{1})|+|Seg(a_{2})|\leq n-1)$ which is also in $p-$th iteration. Now note that, the moment $a_{1}$ meets $a_{2}$, it changes to state Gather2 and starts moving together with $a_{1}$ along with its own pebble until they meet $a_{3}$. When all three of them gathers together at $h_{3}$ there is atleast 3 pebbles at $h_{3}$ (two pebbles carried by $a_{1}$, one pebble carried by $a_{2}$, and one pebble remains for $a_{3}$). In this scenario, they change their state to Coloc. Note that in $p-$th iteration only $a_{1}$ changes to state Gather1 as other agents can not change to state Gather1 from either from Wait2, or, Gather2, or, Coloc. So, for this case the claimed result holds.
Case-II: Let $a_{0}$ was in state Fetch while it was destroyed by the Byzantine black hole node $v_{b}$. In this case, none of the alive agents see more than one pebble at their $home$ after they return there in the $p-$th iteration. This is because, here the the pebbles left by agents $a_{0},a_{1},a_{2}$ and $a_{3}$ has been collected by agents $a_{3},a_{0},a_{1}$ and $a_{2}$, respectively and when they reach their corresponding $home$, the pebble they carried are the only one there. Thus all of them (except $a_{0}$ as it was destroyed during $p-$th iteration in state Fetch) moves to state Wait2 after reaching $home$ at the $p-$th iteration and after the end of the iteration moved to state Forward and starts the $(p+1)-$th iteration. Since $a_{0}$ was destroyed in the previous iteration it remains unavailable to collect the pebble from $h_{1}$, left by $a_{1}$ in the $(p+1)-th$ iteration. Note that here in this iteration both $a_{1}$ and $a_{2}$ collects pebble from $h_{2}$ and $h_{3}$ ( left by $a_{2}$ and $a_{3}$ respectively) and returns back to $h_{1}$ and $h_{2}$ respectively. But $a_{3}$ upon reaching $h_{0}$ does not find any pebble to collect. In this case it returns to $h_{3}$ without any pebble. Note that here only $a_{1}$ sees more than one pebble upon returning back to $h_{1}$ in the $(p+1)-$th iteration. So only $a_{1}$ changes to state Gather1 at round $n+2+|Seg(a_{1})|$. Now by similar argument as in Case-I, the rest follows.
After 3 agents gather at a safe node with at least 3 pebbles they execute the algorithm PerpExplore-Coloc-Pbl which solves the problem of PerpExploration-BBH on a ring $R$ for 3 co-located agents with at least one pebble each agent (refer, Theorem 4.2). Thus this proves Theorem 5.2. which is stated again below.
Statement of Theorem 5.2: Algorithm PerpExplore-Scat-Pbl solves PerpExploration-BBH problem of a ring $R$ with 4 synchronous and scattered agents under the pebble model of communication where each agents are equipped with a pebble.
10.3 Modification of PerpExplore-Scat-Pbl to include the cases where starting positions has multiplicity
@@
In the discussion of our Algorithm 2, we have considered that 4 agents are scattered along 4 distinct nodes (i.e., each node with multiplicity 1). Here we describe how our algorithm (Algorithm 2) works for the remaining cases, (i.e., when 4 agents are scattered among 3 or 2 nodes initially) by slight modification. Observe that, in these remaining cases there exists at least one node, where multiplicity is greater than 1 initially. Let there exists a node with multiplicity $3$, in this case, these agents directly start executing PerpExplore-Coloc-Pbl. While execution of the algorithm, if they encounter the fourth agent somewhere along $R$, they just ignore this agent (note that the IDs, of all the 3 initially co-located agents are already collected by each other) while executing their current algorithm. On the other hand, if initially there exists a node with multiplicity greater than 1 and less than 3, then in that case, only the lowest ID agent (say $a_{1}$) at the current node, changes its state Forward, whereas the other agent at the current node (say $a_{2}$) changes its state Backup-Wait at some round say $t_{0}$ $(>0)$. The agent in state Forward, continue executing the algorithm PerpExplore-Scat-Pbl, whereas the agent in state Backup-Wait, waits at the current node for $12n+$ 1 rounds (i.e., until $t_{0}+2n+1$-th round) and then checks for anomaly in the next round (i.e., if the current node has more pebble than current number of agents). If such anomaly exists after $t_{0}+2n+1$ round, then that implies that $a_{1}$ is already in state Gather1. It is because, $a_{1}$ while it returned back to its $home$ carrying a pebble in state Fetch at round $t_{0}+n+size+2$ round (where $size$ is the length of $Seg(a_{1})$), it encounters the first anomaly, i.e., the number of pebble is more than number agents. In that case, $a_{1}$ directly changes its state to Gather1, and further waits at $home$ for $n-size-1$ rounds, i.e., till $t_{0}+2n+1$ rounds. After which it starts to move clockwise, and on the other hand $a_{2}$ also starts moving clockwise, while changing its state to Gather2 from Backup-Wait. This guarantees that if the anomaly is detected at a multiplicity then both $a_{1}$ and $a_{2}$ moves together to gather with the third agent, which is in state Wait2. Now if the anomaly is not detected at the multiplicity, that is $a_{2}$ at round $t_{0}+2n+1$ finds no anomaly then it moves into state Backup whereas $a_{1}$ is in state Wait2 at the same node. Both of them waits until round $13n+$ 1 and at round $23n+$ 2, $a_{1}$ changes its state to Forward again (refer to Observation 1) and $a_{2}$ changes its state to Backup-Wait again. If some other agent detects anomaly it must meet $a_{1}$ and $a_{2}$ at their $home$ at some round $t$ where $t_{0}+2n+2\leq t\leq t_{0}+3n+1$. In that case they find that there are 3 agents at their $home$ and all of them change to state Coloc and executes algorithm PerpExplore-Coloc-Pbl.
11 Algorithm PerpExplore-Scat-Whitbrd